OpenStudy (anonymous):

hi i need to determine whether the following improper integral is convergent or divergent and determine it's value S b=infinity a=0 (e^x)/(e^e^x) dx

OpenStudy (anonymous):

**the "S" is the integral sign

OpenStudy (anonymous):

**the "S" is the integral sign

OpenStudy (nikvist):

\[y=e^{e^x},\ln y=e^x,\frac{y'}{y}=e^x,\frac{dy}{y}=e^xdx\] \[\int\limits_0^{+\infty}\frac{e^x}{e^{e^x}}dx= \int\limits_e^{+\infty}\frac{dy/y}{y}=\left(-\frac{1}{y}\right)_e^{+\infty}=\frac{1}{e}\]